OpenStudy (anonymous):

WORD PROBLEM HELP..... (use 2 variables to setup & solve) PODCASTS: Steve subscribed to 10 podcasts for a total of 340minutes. He used his two favorite tags, Hobbies & Recreation and Soliloquies. Each of the Hobbies and Recreation episodes lasted about 32 minutes. Each Soliloquies episode lasted 42 minutes. To how many of each tag did Steve subscribe? Let x = H & R podcasts Let y = Soliq. podcasts

OpenStudy (anonymous):

then your two equations are \[x+y=10\] since he subscribes to ten in total. also you have \[32x+42y=340\] because if each of the \(x\) H &R last 32 minutes, then the total number of minutes for them is \(32x\) likewise for the others

OpenStudy (anonymous):

yes ok I'm understanding as you explain and I re-read question

OpenStudy (anonymous):

solve for \(y\) in the first equation by \(y=10-x\) substitute in the second equation to get \[32x+42(10-x)=340\] and solve that one

OpenStudy (anonymous):

32x+420-42x=340 combine like terms. -10x+420=340 ..... then -10x=-80 so x=8

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

yes and then y=2. Thank you for explaining and helping me out. I greatly appreciate it. I really cannot stand word problems. They are difficult for me.

OpenStudy (anonymous):

again thanks & have a good night.